# Ondřej Mokrý Dept. of Telecommunications

### v1

@@ Abstract
The paper focuses on inpainting missing parts of an audio signal spectrogram. First, a recent successful approach based on an untrained neural network is revised and its several modifications are proposed, improving the signal-to-noise ratio of the restored audio. Second, the Janssen algorithm, the autoregression-based state-of-the-art for time-domain audio inpainting, is adapted for the time-frequency setting. This novel method, coined Janssen-TF, is compared to the neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
@@
audio inpainting, autoregression, deep prior, spectrogram.
@@ I Introduction
The time-domain inpainting represents the most natural case from the point of view of practical applications (signal recovery). Nonetheless, many practical algorithms, such as the classic audio coding algorithms (MP3, AAC, etc.), operate on short-time signal spectra. From this perspective, it is valuable to think about inpainting in the time-frequency domain. [^1] In other fields, one can go even further and require inpainting purely in the frequency domain, see for example [14], [15]. In such a case, a portion of time-frequency (TF) coefficients222 typically in the form of a complex-valued matrix are missing, which need to be recovered to allow a later reconstruction of the time-domain signal. Indeed, there have been a few methods developed in the TF context, all of them relying on neural modeling [16, 17, 18]. The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not require training in the common sense.
Note that a compact gap in time bears some fundamental differences compared with a gap in the form of missing a number of consecutive columns of a spectrogram. This follows from the nature of time-frequency transforms, which use windows with overlaps to compute spectrograms, see for instance [19]. In particular, a gap in a spectrogram does not fully correspond to a time-domain gap and vice versa; there is always an amplitude fading in one of the domains.
The goal of this paper is to adopt the Janssen algorithm [2], proven successful in the time-domain case, to the TF-domain case, and compare its performance with the recent deep-prior-based approach of [18], both using objective metrics and by listening tests.
@@ II Deep Prior Audio Inpainting (DPAI)
The deep prior, introduced in [20], is a concept used for the reconstruction of corrupted signals using an “untrained” convolutional neural network (CNN). The signal reconstruction involves only a proper CNN architecture and the corrupted signal. The deep prior concept has also proven that the network architecture matters much more than previously thought.
@@ II-A Application of Deep Prior to Audio Inpainting
Deep prior reconstruction can be utilized in a variety of applications [20]. In the context of the time-domain audio inpainting, consider a corrupted signal $\mathbf{x}^{\textup{cor}}\!\in\mathbb{R}^{N}$, having missing regions (filled with zeros), compared with the original signal $\mathbf{x}^{\textup{orig}}\!\in\mathbb{R}^{N}$. Formally, if a binary mask $\mathbf{m}\in\{0,1\}^{N}$ is utilized to represent the missing sections of the signal, the relation $\mathbf{x}^{\textup{cor}}=\mathbf{m}\odot\mathbf{x}^{\textup{orig}}$ holds, where $\odot$ represents the elementwise product. Such a relation naturally characterizes the feasible set $\mathit{\Gamma}_{\textup{T}}=\{\mathbf{x}\,|\,\mathbf{m}\odot\mathbf{x}=\mathbf{m}\odot\mathbf{x}^{\textup{orig}}\}$.
@@
where $E$ is a differentiable data fidelity function measuring the closeness of its two inputs. The optimal $\theta^{*}$ can be obtained using an iterative optimizer, e.g., Adam [21]. The crucial observation is that the optimization must be interrupted before it reaches $\theta^{*}$, for which we would have $f_{\theta^{*}\!}(\mathbf{n})=\mathbf{x}^{\textup{cor}}$. Instead, early stopping of the training process is desired, resulting in an estimate $\hat{\theta}$ and a corresponding $\hat{\mathbf{x}}=f_{\hat{\theta}}(\mathbf{n})$ [20].
The authors of [18] used an analog approach to inpaint audio in the transformed domain. Through the use of the short-time Fourier transform (STFT), an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{X}\in\mathbb{C}^{T\times F}$. Therefore the CNN operates on spectrograms and $\hat{\mathbf{X}}=f_{\hat{\theta}}(\mathbf{N})$ for a random but fixed matrix $\mathbf{N}$. Once $\hat{\mathbf{X}}$ has been established, the time-domain solution is obtained by the inverse STFT.
Regarding the mask, in [18] only the case of completely missing columns of the spectrogram is considered, i.e., for the TF mask $\mathbf{M}\in\{0,1\}^{T\times F}$ it holds that each column is either full of ones or zeros. Nevertheless, generalization to other TF masks is straightforward. The observed spectrogram $\mathbf{X}^{\textup{cor}}$ is finally modeled as $\mathbf{X}^{\textup{cor}}\!=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}$, where $\mathbf{X}^{\textup{orig}}$ corresponds to $\mathbf{x}^{\textup{orig}}$ in the time domain. Note that by analogy with the time-domain case, we have the feasible set $\mathit{\Gamma}_{\textup{TF}}=\{\mathbf{X}\,|\,\mathbf{M}\odot\mathbf{X}=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}\}$.
@@
where $\mathrm{MSS}$ is the multi-scale spectrogram loss [22]. In [18], the weights $\alpha_{1}$ and $\alpha_{2}$ have values 1 and 0.1 respectively.
@@ II-B Architecture
II-C Modifications
The DPAI reconstructs the whole $\hat{\mathbf{X}}$. Yet, considering that only a part of the spectrogram is not known, the authors of [18] evaluate the performance of their method after substituting the coefficients of $\hat{\mathbf{X}}$ that correspond to ones in the mask $\mathbf{M}$ with the observed data. Formally, such an action is a projection
@@ II-C Modifications
onto the set $\mathit{\Gamma}_{\textup{TF}}$. Note that a resulting sharp transition between the regions is smeared in the time domain due to the nature of the STFT [19]. We call the modified reconstructions with context, while the others using the entire $\hat{\mathbf{X}}$ are without context.
The other modifications come from the fact that the reconstruction always starts from a random initialization of both $\theta$ and $\mathbf{N}$. As a consequence, the objective and the subjective qualities of the reconstructed signal might slightly vary, based on the particular realization. We observed that averaging the reconstructions from multiple independent reconstructions of the same signal improves the objective performance metrics. Obviously, the cost of such an approach grows linearly with the number of trials. We also noticed that a slight improvement can be assured by averaging a number of final epochs of an individual reconstruction.
In the experimental part, we will study whether the proposed modification of DPAI brings a significant performance improvement.
@@ III Janssen spectrogram inpainting (Janssen-TF)
In the inpainting method proposed by Janssen et al. [2], the restored audio signal is supposed to follow the AR model (4) while both the missing samples in the time-domain and the unknown AR parameters $\mathbf{a}$ are being estimated. Janssen et al. approach the problem iteratively by repeating two principal steps in the $i$-th iteration:
@@
Step 1 can be solved efficiently using standard methods, such as the Durbin–Levinson recursion (see e.g. the function lpc in Matlab). Regarding step 2, we propose to adapt the Janssen iteration to the TF setting, denoted Janssen-TF. In our adaptation, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, while the signal follows the estimated AR model. To formalize this modification, note that the error in (4) can be written as $\mathbf{e}=\mathbf{A}\mathbf{x}$, where $\mathbf{A}\in\mathbb{R}^{(N+p)\times N}$ is a Toeplitz matrix composed using the (fixed) coefficients $\mathbf{a}$. We can then pose the proposed modification of step 2 as
@@
where $\iota_{\mathit{\Gamma}_{\textup{TF}}}$ denotes the indicator function of the convex set $\mathit{\Gamma}_{\textup{TF}}$ of feasible spectrograms. [^3] Note that the original Janssen algorithm minimizes the function $\frac{1}{2}\|\mathbf{A}\mathbf{x}\|+\iota_{\mathit{\Gamma}_{\textup{T}}}(\mathbf{x})$ in subproblem 2. As such, the problem (5) is convex and it is solvable using standard numerical methods. We propose to employ the alternating direction method of multipliers (ADMM), which includes two principal steps [26, Sec. 3.1.1]:
@@ IV Experiments
IV-A Setup and metrics
For our experiments, we adopt the audio dataset [^5] https://github.com/fmiotello/dpai/ used for the evaluation of DPAI [18]. It consists of 6 music and 2 speech recordings of 5 seconds in length, sampled at 16 kHz. Only minor modifications were made to the original distribution of gaps (i.e., the placement of the missing spectrogram columns), such that each test excerpt includes 5 gaps of the same length. The gap lengths range from 1 up to 6 missing columns. In turn, the entire test set comprises of $8\times 6=48$ corrupted audio signals. Even though the dataset is relatively small and includes only a few types of audio signals, we decided to stick to it because its use justifies a fair comparison of the proposed method with DPAI. The choice allows the use of exactly the same U-net architecture of DPAI from [18]. In particular, we choose their “best2” setting. The number of epochs of the Adam optimizer is 5000.
The STFT (i.e., the operator $\mathbf{L}$) utilizes the 2048-sample-long Hann window (which corresponds to 128 ms), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^6] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
To objectively evaluate the reconstruction quality, we assess the difference between the inpainted signal $\hat{\mathbf{x}}$ and a clean reference $\mathbf{x}^{\textup{orig}}$ using two standard metrics:
@@ IV-A Setup and metrics
    a subjective listening test containing a limited number of excerpts. The MUSHRA-type test [32] was performed that includes the reference signal, the DPAI reconstruction with context, the DPAI reconstruction without context, the DPAI with context but averaged result from four random initializations, the gapwise Janssen method in the time domain [12], the Janssen-TF-ADMM, and finally the hidden anchor, i.e., the signal reconstructed from the spectrogram with gaps.
The participants were asked to score the perceived similarity of the signals with the reference on a continuous scale from 0 to 100, with the help of the webMUSHRA environment [^7] https://github.com/audiolabs/webMUSHRA/. The test was run in a quiet music studio, using a professional sound card and headphones. The listening conditions were identical for all the participants. Following the recommendation [32], 11 assessors out of 18 were selected through post-screening.
@@ IV-B Results
First, we evaluate the proposed modifications of DPAI. As seen in Fig. 1, averaging the realizations brings improvement in quality in terms of SNR, but it comes at the cost of a higher time complexity. The differences observed in Fig. 1 were further tested using the Wilcoxon signed rank test [^8] https://www.mathworks.com/help/stats/signrank.html, which proved that the median of the SNR using DPAI with averaging is significantly higher compared to the original DPAI with context. On the other hand, a significant difference in case of the ODG metric has not been proven.
001010101020202020303030304040404050505050001010101020202020303030304040404050505050SNR (dB), DPAI with contextSNR (dB), DPAI averagedGaplength1 2 3 4 5 6 44-4- 433-3- 322-2- 211-1- 10044-4- 433-3- 322-2- 211-1- 100ODG, DPAI with contextODG, DPAI averaged
Figure 1: Comparison of DPAI reconstruction with context versus the proposed DPAI variant averaging four random initializations.
In contrast to the averaging of realizations, the averaging across the final epochs from a single realization does not increase the computational load and brings better quality mostly to audio recordings reconstructed without context—the greatest improvement (averaged over test signals) was 0.16 ODG for the gap of 2 columns and 15 epochs averaged. However, this strategy has a negligible impact when considering reconstructions with context.
The objective comparison that includes the proposed Janssen-TF method is shown in Fig. 2. The first observation is that the time-domain-based Janssen method [12] cannot compete with the other methods. This is due to the incompatibility of gaps in the time and time-frequency domains (a sharp mask in the spectrogram domain affects a longer signal portion when transformed into the time domain). While the average performance of the DPAI variants mostly lie within the confidence intervals of each other, suggesting limited significance of the differences, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except for the longest gaps).
Figure 2: Comparison of the inpainting methods using objective metrics – SNR (top) and ODG (bottom). The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [33, p. 160]. If the intervals do not overlap, it may be concluded that the difference of the means is statistically significant.
To support the objective evaluation, Fig. 3 presents the results of the subjective listening test. Its results correlate well with the objective evaluation; in particular, Janssen-TF-ADMM is superior to the competing methods. Moreover, this result is statistically significant in terms of medians. On the other hand, the DPAI variants scored almost the same. This is in line with the ODG results in Fig. 1, even though the listening test featured only a selection of the test signals.
anchorreferenceDPAI without contextDPAI with contextDPAI averagedgapwise JanssenJanssen-TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score
@@
A single run of DPAI takes ca 19 minutes on the NVIDIA Tesla V100S GPU with 32 GB of memory, regardless the number of missing spectrogram columns. Computing a single signal with Janssen-TF-ADMM takes 10 to 20 minutes, directly proportional to the number of columns; this corresponds to a PC with Intel Core i7 3.40 GHz processor, 32 GB RAM.

### v2

@@ Abstract
The paper focuses on inpainting missing parts of an audio signal spectrogram. The autoregression-based Janssen algorithm, the state-of-the-art for the time-domain audio inpainting, is adapted for the time-frequency setting. This novel method, termed Janssen-TF, is compared to the deep-prior neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
@@
audio inpainting, autoregression, deep prior, time-frequency, spectrogram.
@@ I Introduction
The time-domain inpainting represents the most natural case from the point of view of the application to signal recovery. Nonetheless, many practical algorithms, such as classic audio codecs (MP3, AAC, etc.), operate on short-time signal spectra. From this perspective, it is valuable to think about inpainting in the time-frequency domain. [^1] In other fields, one can go even further and require inpainting purely in the frequency domain, see for example [14], [15]. In such a case, a portion of time-frequency (TF) coefficients222 typically in the form of a complex-valued matrix are missing, which need to be recovered to allow a later reconstruction of the time-domain signal. Indeed, there have been a few methods developed in the TF context, all of them relying on neural modeling [16, 17, 18]. The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not need training in the most common sense.
Note that a compact gap in time bears some fundamental differences compared with a gap in the form of a missing number of consecutive columns of a spectrogram. This follows from the nature of time-frequency transforms, which use windows with overlaps to compute spectrograms, see for instance [19]. In particular, a gap in a spectrogram does not fully correspond to a time-domain gap and vice versa; there is always an amplitude fading in one of the domains thanks to the windowing.
The goal of this paper is to adapt the Janssen algorithm [2], proven successful in the time-domain case, for the TF-domain case, and compare its performance with the recent deep-prior-based approach of [18], both using objective metrics and by listening tests.
@@ II Deep Prior Audio Inpainting (DPAI)
The deep prior, introduced in [20], is a concept used for the reconstruction of corrupted signals using a convolutional neural network (CNN). In contrast to most CNN applications, [20] does not utilize a training dataset; the reconstruction only involves a proper CNN architecture and the corrupted signal. The deep prior concept has also proven that the network architecture matters much more than previously thought.
@@ II-A Application of Deep Prior to Audio Inpainting
The deep prior reconstruction can be utilized in a variety of applications, as outlined in [20]. In the context of the time-domain audio inpainting, consider a corrupted signal $\mathbf{x}^{\textup{cor}}\!\in\mathbb{R}^{N}$, having missing regions (filled with zeros), compared with the original signal $\mathbf{x}^{\textup{orig}}\!\in\mathbb{R}^{N}$. Formally, if a binary mask $\mathbf{m}\in\{0,1\}^{N}$ is utilized to represent the missing sections of the signal, the relation $\mathbf{x}^{\textup{cor}}=\mathbf{m}\odot\mathbf{x}^{\textup{orig}}$ holds, where $\odot$ represents the elementwise product. Such a relation naturally characterizes the feasible set $\mathit{\Gamma}_{\textup{T}}=\{\mathbf{x}\,|\,\mathbf{m}\odot\mathbf{x}=\mathbf{m}\odot\mathbf{x}^{\textup{orig}}\}$.
@@
where $E$ is a differentiable data fidelity function measuring the closeness of its two inputs. The optimal $\theta^{*}$ can be obtained using an iterative optimizer, e.g., Adam [21]. The crucial observation is that the optimization must be interrupted before it reaches $\theta^{*}$, for which we would have $f_{\theta^{*}\!}(\mathbf{n})=\mathbf{x}^{\textup{cor}}$, meaning basically overfitting. Instead, early stopping of the training is desired, resulting in an estimate $\hat{\theta}$ and a corresponding $\hat{\mathbf{x}}=f_{\hat{\theta}}(\mathbf{n})$ [20].
The authors of [18] used an analog approach to inpaint audio in the transformed domain. Through the use of the short-time Fourier transform (STFT), denoted $\mathbf{L}$ in the following, an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{L}\mathbf{x}=\mathbf{X}\in\mathbb{C}^{T\times F}$. Therefore the CNN operates on spectrograms and $\hat{\mathbf{X}}=f_{\hat{\theta}}(\mathbf{N})$ for a random but fixed matrix $\mathbf{N}$. Once $\hat{\mathbf{X}}$ has been established, the time-domain solution is obtained by the inverse STFT.
Regarding the mask, in [18] only the case of completely missing columns of the spectrogram is considered, i.e., for the TF mask $\mathbf{M}\in\{0,1\}^{T\times F}$ it holds that each column is either full of ones or zeros. Nevertheless, a generalization to other TF masks is straightforward. The observed spectrogram $\mathbf{X}^{\textup{cor}}$ is finally modeled as $\mathbf{X}^{\textup{cor}}\!=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}$, where $\mathbf{X}^{\textup{orig}}$ corresponds to $\mathbf{x}^{\textup{orig}}$ in the time domain, i.e., $\mathbf{X}^{\textup{orig}}=\mathbf{L}\mathbf{x}^{\textup{orig}}$. Note that by analogy with the time-domain case, we have the feasible set $\mathit{\Gamma}_{\textup{TF}}=\{\mathbf{X}\,|\,\mathbf{M}\odot\mathbf{X}=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}\}$.
@@
where MSE is the Mean Squared Error and $\mathrm{MSS}$ is the multi-scale spectrogram loss [22]. In [18], the weights $\alpha_{1}$ and $\alpha_{2}$ have values 1 and 0.1 respectively.
@@ II-B Architecture
II-C Variants
The DPAI reconstructs the whole $\hat{\mathbf{X}}$. Yet, considering that only a part of the spectrogram was unavailable, the authors of [18] evaluate the performance of their method after injecting the reliable coefficients back to the solution as a postprocessing step. Formally, such an action is a projection
@@ II-C Variants
onto the set $\mathit{\Gamma}_{\textup{TF}}$. Note that a resulting sharp transition between the regions is smeared in the time domain due to the nature of the STFT [19]. We call the reconstructions modified by (3) with context, while the others using the entire $\hat{\mathbf{X}}$ are without context. In the experimental part, we will study whether such a modification brings a significant performance improvement.
@@ III Janssen spectrogram inpainting (Janssen-TF)
In the time-domain inpainting method proposed by Janssen et al. [2], the restored audio signal is supposed to follow the AR model (4) while both the missing samples in the time-domain and the unknown AR parameters $\mathbf{a}$ are being estimated. Janssen et al. approach the problem iteratively by performing two principal steps in the $i$-th iteration:
@@
Step 1 can be solved efficiently using standard methods, such as the Durbin–Levinson recursion (see e.g. the function lpc in Matlab). Regarding step 2, we propose to adapt the Janssen iteration to the TF setting, denoted Janssen-TF. In our adaptation, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, while the signal itself follows the estimated AR model. To formalize this modification, note that the error in (4) can be written as $\mathbf{e}=\mathbf{A}\mathbf{x}$, where $\mathbf{A}\in\mathbb{R}^{(N+p)\times N}$ is a Toeplitz matrix composed using the (fixed) coefficients $\mathbf{a}$. We can then pose the proposed modification of step 2 as
@@
where $\mathbf{L}$ is the STFT and $\iota_{\mathit{\Gamma}_{\textup{TF}}}$ denotes the indicator function of the convex set $\mathit{\Gamma}_{\textup{TF}}$ of feasible spectrograms. [^3] Note that the original Janssen algorithm minimizes the function $\frac{1}{2}\|\mathbf{A}\mathbf{x}\|+\iota_{\mathit{\Gamma}_{\textup{T}}}(\mathbf{x})$ in subproblem 2. As such, the problem (5) is convex and it is solvable using standard numerical methods. We propose to employ the alternating direction method of multipliers (ADMM), which includes two principal steps [26, Sec. 3.1.1]:
@@ IV Experiments
IV-A Data, setup and metrics
For our first experiment, we adopt the audio dataset [^5] https://github.com/fmiotello/dpai/ used for the evaluation of DPAI [18]. It consists of 6 music and 2 speech recordings of 5 seconds in length, sampled at 16 kHz. Only minor modifications were made to the original distribution of gaps (i.e., the placement of the missing spectrogram columns), such that each test excerpt includes 5 gaps of the same length. The gap lengths range from 1 to 6 missing columns. In turn, the entire test set comprises $8\times 6=48$ corrupt spectrograms. Even though the dataset is relatively small and includes only a few types of audio signals, it provides a fair comparison of the proposed method with DPAI—the choice allows the use of exactly the same U-net architecture of DPAI from [18]. In particular, we chose their “best2” setting, and 5000 epochs of the Adam optimizer.
To validate the results on a larger dataset, we hand-picked 60 music recordings of various complexity from the IRMAS dataset [31, 32]. To enable the use of DPAI, the recordings were subsampled to 16 kHz and shortened to 5 seconds.
The STFT (i.e., the operator $\mathbf{L}$) utilizes a 2048-sample-long Hann window (which corresponds to 128 ms), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^6] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
We compare the proposed method Janssen-TF-ADMM with three baseline methods:
  • DPAI reconstruction with context,
  • DPAI reconstruction without context,
  • gapwise Janssen method in the time domain [12].
The gapwise method is a twist to the original Janssen algorithm [2], which, instead of processing by means of overlapping frames, makes use of a single segment consisting of the whole gap and its nearby context.
To objectively evaluate the reconstruction quality, we assess the difference between the inpainted signal $\hat{\mathbf{x}}$ and the clean reference $\mathbf{x}^{\textup{orig}}$ using two standard metrics:
@@ IV-A Data, setup and metrics
    a subjective listening test containing a limited number of excerpts. The MUSHRA-type test [34] was performed, which includes the reference signal, the competing recosntructions, and finally the hidden anchor, i.e., the signal reconstructed from the spectrogram with gaps. The participants were asked to score the perceived similarity of the signals with the reference signal on a continuous scale from 0 to 100, with the help of the webMUSHRA environment [^7] https://github.com/audiolabs/webMUSHRA/. The test was run in a quiet music studio, using a professional sound card and headphones. The listening conditions were identical for all the participants. Following the recommendation [34], 11 assessors out of 18 were selected through post-screening.
@@ IV-B Results
The objective comparison, including the proposed Janssen-TF method, is shown in Fig. 1 for the case of the DPAI dataset. The first observation is that the time-domain-based Janssen method [12] is not competitive with the other approaches. This can be explained by the incompatibility of gaps in the time and time-frequency domains (a sharp mask in the spectrogram domain affects a longer signal portion when transformed into the time domain, and the whole affected segment is treated as missing). While the average performance of the DPAI variants mostly lies within the confidence intervals of each other, suggesting limited significance of the differences, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except the longest gaps).
Figure 1: Comparison of the inpainting methods using objective metrics – SNR (top) and ODG (bottom). The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [35, p. 160]. If the intervals do not overlap, it may be concluded that the difference of the means is statistically significant.
To support the objective evaluation on the first dataset, Fig. 2 presents the results of the subjective listening test. Its results correlate well with the objective evaluation; in particular, Janssen-TF-ADMM is superior to the other methods. Moreover, this result is statistically significant in terms of medians. On the other hand, the DPAI variants scored almost the same.
anchorreferenceDPAI without contextDPAI with contextgapwise JanssenJanssen--TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score
@@
Finally, Fig. 3 shows the evaluation on the IRMAS subset comprising 60 signals. This evaluation supports the previous analysis and leads to the conclusion that the proposed method is superior to the others. For shorter gaps, DPAI with context is preferable over DPAI not exploiting context.
Figure 3: SNR and ODG results on the IRMAS dataset.
Regarding computational demands, a single run of DPAI takes ca 19 minutes on the NVIDIA Tesla V100S GPU with 32 GB of memory, regardless of the number of missing spectrogram columns. Reconstructing a single signal with Janssen-TF-ADMM takes 10 to 20 minutes, proportional to the number of columns; this performance corresponds to a PC with Intel Core i7 3.40 GHz processor, 32 GB RAM.